# Weak Existence and Uniqueness for Super-Brownian Motion with Irregular Drift

### v1

@@ 6.3 Proof of Approximate Duality
In this section, we assume that there exists a weak solution $(X_{t})_{t\geq 0}$ to (1.1) with $X_{0}\sim\eta\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ such that (1.5) is satisfied and prove an (approximate) duality relation. When $h$ is continuous, the existence of a solution was already resolved in [25, Theorem 1.1]. The existence of solutions when $h$ is discontinuous is proven in Section 7 and relies on the duality relation for the continuous case. In order to establish the duality relation, we consider approximation of the process $Y$ defined in (6.5). To this end, we denote by $Y^{n}$ the process constructed in (6.5) with $\nu$ replaced by
@@
Note that
$\begin{split}&\mathbb{E}[X_{t}(x)]=\mathbb{E}[\big{(}S_{t}X_{0}\big{)}(x)]+\int_{0}^{t}\int_{\mathbb{R}}p_{t-s}\mathbb{E}[h(X_{s}(y))]dyds\\ &\leq\mathbb{E}[\big{(}S_{t}X_{0}\big{)}(x)]+C\int_{0}^{t}\int_{\mathbb{R}}p_{t-s}(x-y)dyds,\end{split}$
@@
The proof follows along similar lines as the proof of [21, Lemma 2.6], but we include it for the sake of completeness. Let $(\mu_{k})_{k\geq 1}$ be a sequence in $\mathcal{C}_{c}^{\infty}$ that converges weakly to $\mu$. Then it follows that $V_{\cdot}(\mu_{k})\in\mathcal{C}^{1,2}(\mathbb{R}_{+}\times\mathbb{R})\cap L^{2}_{loc}(\mathbb{R}_{+}\times\mathbb{R})$, where $\psi\in L^{2}_{loc}(\mathbb{R}_{+}\times\mathbb{R})$ if
@@
$\mathbb{E}_{X}\Big{[}\int_{t}^{T}\exp(-\langle V_{T-s}(\mu_{k}),X_{s}\rangle+v_{T-s})dM_{s}(V_{T-\cdots}(\mu_{k}))\Big{]}=0.$ (6.52)
@@
shows that $\int_{0}^{t}\exp(-\langle V_{T-s}(\mu_{k}),X_{s}\rangle+v_{T-s})dM_{s}(V_{T-s}(\mu_{k}))$ is a martingale, which implies (6.52). It remains to show the convergence as $k\rightarrow\infty$. Using [22, Lemma 2.1], the dominated convergence theorem and setting $\tilde{t}=T-t$ in (6.51) yields
@@
Let $(X_{t})_{t\geq 0}$ be a solution to (1.1) with $X_{0}\sim\eta\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ such that (1.5) holds and let $a$ be defined as in (A). Let $Y^{n}$ be as above, independent of $X$. Then there exists a time $T_{0}>0$ only depending on $\langle\mu,1\rangle$ and $a$ such that for all $T\leq T_{0}$ we have
@@
First, choose $T_{0}>0$ small enough such that Lemma 6.5 is satisfied with $\gamma=2a$ and let $T\leq T_{0}$. Furthermore, set $v_{t}=-a\int_{0}^{t}\langle Y^{n}_{s},1\rangle ds$ for $t\geq 0$. By Lemma 6.12 we have
@@
since $t\mapsto X_{t}$ is continuous and $J_{\bar{T}_{n,k+1}-}=J_{s}$ for all $s\in(\bar{T}_{n,k},\bar{T}_{n,k+1})$. Let $\boldsymbol{\Delta}Y_{s}=Y_{s}-Y_{s-}$ for all $s\geq 0$, then we get
$\begin{split}&\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k+1}}}\exp\Big{(}-\langle Y_{\bar{T}_{n,k+1}}^{n},X_{T-\bar{T}_{n,k+1}}\rangle-a\int_{0}^{\bar{T}_{n,k+1-}}\langle Y_{s}^{n},1\rangle ds\Big{)}\Big{]}\\ &=\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\Big{(}-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle+v_{\bar{T}_{n,k+1}}\Big{)}\Big{]}\\ &+\exp\big{(}v_{\bar{T}_{n,k+1}}\big{)}\mathbb{E}_{X}\Big{[}(-1)^{J_{\bar{T}_{n,k+1}}}\exp\big{(}-\langle Y_{\bar{T}_{n,k+1}}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big{)}\\ &\hskip 199.16928pt-(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\big{(}-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big{)}\Big{]}\\ &=\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k}}}\exp\Big{(}-\langle Y_{\bar{T}_{n,k}},X_{T-\bar{T}_{n,k}}\rangle+v_{\bar{T}_{n,k}}\Big{)}\Big{]}\\ &+\mathbb{E}_{X}\Big{[}\int_{\bar{T}_{n,k}}^{\bar{T}_{n,k+1}}(-1)^{J^{n}_{s-}}\exp\Big{(}-\langle Y_{s-}^{n},X_{T-s}\rangle+v_{s}\Big{)}\langle Y_{s-}^{n},h(X_{T-s})-a\rangle ds\Big{]}\\ &-\exp\big{(}v_{\bar{T}_{n,k+1}}\big{)}\mathbb{E}_{X}\Big{[}(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\big{(}-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big{)}\\ &\times\Big{(}(1-\exp(-\langle\boldsymbol{\Delta}Y^{n}_{\bar{T}_{n,k+1}},X_{T-\bar{T}_{n,k+1}}))\mathbbm{1}_{M_{n,k}=1}+\big{(}1+\exp\big{(}-\langle\boldsymbol{\Delta}Y^{n}_{\bar{T}_{n,k+1}},X_{T-\bar{T}_{n,k+1}}\rangle\big{)}\big{)}\mathbbm{1}_{M_{n,k}=2}\Big{)}\Big{]},\end{split}$
@@
By Lemma 6.14 we have
@@ 7 Proof of Weak Existence
Note, that we only need to proof weak existence in the case where $h$ is discontinuous due to [25, Theorem 1.1]. That means we need to consider the case where $b_{0}\neq b_{1}\neq 0$. For simplicity, we thus only consider the case where $\nu^{1}=\nu^{2}\equiv 0$. The prove relies on considering a sequence $(X^{n})_{n\geq 1}$ of solutions to (1.1), where $h$ is approximated by continuous functions $h_{n}$ given by
@@
where $d_{1},d_{2}$ and $a$ are given according to (A). The sequence of functions $(h_{n})_{n\geq 1}$ converges boundedly pointwise to $h$. As mentioned above, we get the existence of a weak solution $X^{n}$ by applying [25, Theorem 1.1] for each $n\in N$. When $h$ is replaced by $h_{n}$, we also get the weak uniqueness of solutions by applying the results from Section 8. Furthermore, we can apply the duality relation from Corollary 6.15 for $X^{n}$, which is conditional on existence.
@@ 7.1 Tightness
The proof follows from standard estimates as in the proof of [25, Theorem 2.6]. ∎
Lemma 7.1 demonstrates by [25, Lemma 6.3] that the sequence $(X^{n})_{n\geq 1}$ of $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}(\mathbb{R}))$ valued random variables is tight. Thus, by Prohorov’s theorem and Skorohod’s representation theorem, there exists a sub-sequence $(X^{n_{k}})_{k\geq 1}$ and another sequence $(\tilde{X}^{n_{k}})_{k\geq 1}$ of $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}(\mathbb{R}))$ valued random variables such that the following holds.
@@ 7.2 Convergence of $h_{n}(X^{n})$
The proof is similar to the proof of [3, Lemma 5.3]. First, we prove the $L^{1}$ convergence. Let $T_{0}>0$ be such that Proposition 4.3 holds for $X^{n}$, for all $n\geq 0$. Note that it is possible to choose $T_{0}>0$ uniformly in $n$, since it only depends on $\sup_{n}\nu_{n}^{1}(0,\infty)+\nu_{n}^{2}(0,\infty)$. For simplicity denote $X_{t}(x)$ by $Z$ and $X^{n}_{t}(x)$ by $Z_{n}$. Below we use the fact that almost surely $Z_{n},Z\geq 0$. Thus we get
@@ 8 Proof of Weak Uniqueness
Since $f$ was arbitrary it follows by [13, Corollary 2.3] that $X^{1}_{t}\overset{d}{=}X_{t}^{2}$. ∎
@@
It is enough to prove weak uniqueness on the interval $0,T_{0}]$ and that $\mathbb{P}\circ(X_{T_{0}})^{-1}\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ and that $X_{T_{0}}$ satisfies (1.5), since $T_{0}$ does not depend on the initial value $X_{0}$. This follows along similar lines as Lemma 7.1 and by Lemma 6.11. Thus, it remains to prove uniqueness on $0,T_{0}]$. However, by using Lemma 8.1, this follows along the same lines as the proof of [11, Theorem 4.2]. It is only necessary to show that t $\mathbb{P}\circ(X_{t})^{-1}\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ and that $X_{t}$ satisfies (1.5) for all $t\in[0,T_{0}]$. However, this once again follows along similar lines as Lemma 7.1 and from Lemma 6.11. ∎
@@
Let $T_{0}>0$ be small enough such that
$\mathbb{E}[\operatorname{e}^{|I_{T_{0}}(Y_{0})||}]<\infty$ (9.1)
with $Y_{0}=\infty\delta_{0}$. Such $T_{0}>0$ exists due to Lemma 6.2. For $s=0$ the assertion follows by assumption. Thus, let $s\in(0,T_{0}]$. Using Proposition 4.3 and Remark 4.4 it follows that
@@
where $\bar{T}_{n,k}^{x}$ is time of the $k$-th jump of the process $Y(n\delta_{x})$. That is, $K^{n}_{t}$ is the number of jumps of $Y(n\delta)$ on the time interval $(0,t]$. We furthermore denote $T_{n,k}^{x}=\bar{T}_{n,k}^{x}-\bar{T}_{n,k-1}^{x}$, for $k\geq 1$. We obtain the following decomposition
@@
Finally, we estimate $I_{2}^{n}(s,x)$. To this end, we brake up the event into several parts that we can handle more easily.
@@
$\begin{split}&\mathbb{P}(U_{n,j}^{x}>R_{x}+R,j\in\{1,\ldots k-1\},U_{n,k}^{x}\leq R_{x}+R,K_{s}^{n}=m)\\ &\leq\mathbb{P}(U_{n,k}^{x}\leq R_{x}+R,U_{n,j}^{x}>x-j\eta_{x}\text{ for all }j\in\{1,\ldots k-1\},K_{s}^{n}=m)\\ &\qquad+\mathbb{P}(U_{n,j}^{x}\leq x-j\eta_{x}\text{ for some }j\in\{1,\ldots k-1\},K_{s}^{n}=m)\end{split}$ (9.9)
@@
where $\mathcal{G}_{n,x,k}$ is the $\sigma$-algebra generated by $(U_{n,1}^{x},\ldots U_{n,k-1}^{x},T_{n,1}^{x},\ldots T_{n,k}^{x})$ . For $y\in\mathbb{R}$ we get using Lemma 3.3, Proposition 3.6 and the construction of the process $Y(n\delta_{x})$ that
@@
where we used that $\eta_{x}\leq R_{x}$ for $x\geq R^{\ast}$. This proves the assertion. ∎

### v2

@@ 6.3 Proof of Approximate Duality
In this section, we assume that there exists a weak solution $(X_{t})_{t\geq 0}$ to (1.1) with $X_{0}\sim\eta\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ such that (1.5) is satisfied and prove an (approximate) duality relation. When $h$ is continuous, the existence of a solution was already resolved in [35, Theorem 1.1]. The existence of solutions when $h$ is discontinuous is proven in Section 7 and relies on the duality relation for the continuous case. In order to establish the duality relation, we consider approximation of the process $Y$ defined in (6.6). To this end, we denote by $Y^{n}$ the process constructed in (6.6) with $\nu$ replaced by
@@
Note that, since $h$ is bounded,
$\begin{split}&\mathbb{E}[X_{t}(x)]=\mathbb{E}[\big(S_{t}X_{0}\big)(x)]+\int_{0}^{t}\int_{\mathbb{R}}p_{t-s}\mathbb{E}[h(X_{s}(y))]dyds\\ &\leq\mathbb{E}[\big(S_{t}X_{0}\big)(x)]+C\int_{0}^{t}\int_{\mathbb{R}}p_{t-s}(x-y)dyds\\ &=\mathbb{E}[\big(S_{t}X_{0}\big)(x)]+Ct,\end{split}$
@@
The proof follows along similar lines as the proof of [30, Lemma 2.6], but we include it for the sake of completeness. Let $(\mu_{k})_{k\geq 1}$ be a sequence in $\mathcal{C}_{c}^{\infty}$ that converges weakly to $\mu$. Then it follows that $V_{\cdot}(\mu_{k})\in\mathcal{C}^{1,2}(\mathbb{R}_{+}\times\mathbb{R})\cap L^{2}_{loc}(\mathbb{R}_{+}\times\mathbb{R})$, where $\psi\in L^{2}_{loc}(\mathbb{R}_{+}\times\mathbb{R})$ if
@@
$\mathbb{E}_{X}\Big[\int_{t}^{T}\exp(-\langle V_{T-s}(\mu_{k}),X_{s}\rangle+v_{T-s})dM_{s}(V_{T-\cdot}(\mu_{k}))\Big]=0.$ (6.58)
@@
shows that $\int_{0}^{t}\exp(-\langle V_{T-s}(\mu_{k}),X_{s}\rangle+v_{T-s})dM_{s}(V_{T-s}(\mu_{k}))$ is a martingale, which implies (6.58). It remains to show the convergence as $k\rightarrow\infty$. Using [31, Lemma 2.1], the dominated convergence theorem and setting $\tilde{t}=T-t$ in (6.57) yields
@@
Let $(X_{t})_{t\geq 0}$ be a solution to (1.1) with $X_{0}\sim\eta\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$, $p>2$, such that (1.5) holds and let $a$ be defined as in (A). Let $Y^{n}$ be as above, independent of $X$. Then there exists a time $T_{0}>0$ only depending on $\langle\mu,1\rangle$ and $a$ such that for all $T\leq T_{0}$ we have
@@
First, choose $T_{0}>0$ small enough such that Lemma 6.5 is satisfied with $\gamma=2a$ and $K=\nu((0,\infty])$ and let $T\leq T_{0}$. Furthermore, set $v_{t}=-a\int_{0}^{t}\langle Y^{n}_{s},1\rangle ds$ for $t\geq 0$. Let $k\in\mathbb{N}$ be such that $\bar{T}_{n,k+1}\leq T$. Then, by Lemma 6.12 we have
@@
since $t\mapsto X_{t}$ is continuous and $J_{\bar{T}_{n,k+1}-}=J_{s}$ for all $s\in[\bar{T}_{n,k},\bar{T}_{n,k+1})$. Let $\boldsymbol{\Delta}Y_{s}=Y_{s}-Y_{s-}$ for all $s\geq 0$, then we get
$\begin{split}&\mathbb{E}_{X}\Big[(-1)^{J^{n}_{\bar{T}_{n,k+1}}}\exp\Big(-\langle Y_{\bar{T}_{n,k+1}}^{n},X_{T-\bar{T}_{n,k+1}}\rangle-a\int_{0}^{\bar{T}_{n,k+1}}\langle Y_{s}^{n},1\rangle ds\Big)\Big]\\ &=\mathbb{E}_{X}\Big[(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\Big(-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle+v_{\bar{T}_{n,k+1}}\Big)\Big]\\ &+\exp\big(v_{\bar{T}_{n,k+1}}\big)\mathbb{E}_{X}\Big[(-1)^{J_{\bar{T}_{n,k+1}}}\exp\big(-\langle Y_{\bar{T}_{n,k+1}}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big)\\ &\hskip 199.16928pt-(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\big(-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big)\Big]\\ &=\mathbb{E}_{X}\Big[(-1)^{J^{n}_{\bar{T}_{n,k}}}\exp\Big(-\langle Y_{\bar{T}_{n,k}},X_{T-\bar{T}_{n,k}}\rangle+v_{\bar{T}_{n,k}}\Big)\Big]\\ &+\mathbb{E}_{X}\Big[\int_{\bar{T}_{n,k}}^{\bar{T}_{n,k+1}}(-1)^{J^{n}_{s-}}\exp\Big(-\langle Y_{s-}^{n},X_{T-s}\rangle+v_{s}\Big)\langle Y_{s-}^{n},h(X_{T-s})-a\rangle ds\Big]\\ &-\exp\big(v_{\bar{T}_{n,k+1}}\big)\mathbb{E}_{X}\Big[(-1)^{J^{n}_{\bar{T}_{n,k+1}-}}\exp\big(-\langle Y_{\bar{T}_{n,k+1}-}^{n},X_{T-\bar{T}_{n,k+1}}\rangle\big)\\ &\times\Big((1-\exp(-\langle\boldsymbol{\Delta}Y^{n}_{\bar{T}_{n,k+1}},X_{T-\bar{T}_{n,k+1}}))\mathbbm{1}_{M_{n,k}=1}+\big(1+\exp\big(-\langle\boldsymbol{\Delta}Y^{n}_{\bar{T}_{n,k+1}},X_{T-\bar{T}_{n,k+1}}\rangle\big)\big)\mathbbm{1}_{M_{n,k}=2}\Big)\Big],\end{split}$
@@
By Lemma 6.14 there exists $T_{0}>0$ such that for all $T\leq T_{0}$ we have
@@ 7 Proof of Weak Existence
Note, that we only need to proof weak existence in the case where $h$ is discontinuous due to [35, Theorem 1.1]. That means we need to consider the case where $b_{0}\neq b_{1}\neq 0$. For simplicity, we thus only consider the case where $\nu^{1}=\nu^{2}\equiv 0$. The prove relies on considering a sequence $(X^{n})_{n\geq 1}$ of solutions to (1.1), where $h$ is approximated by continuous functions $h_{n}$ given by
@@
where $d_{1},d_{2}$ and $a$ are given according to (A). The sequence of functions $(h_{n})_{n\geq 1}$ converges boundedly pointwise to $h$. As mentioned above, we get the existence of a weak solution $X^{n}$ by applying [35, Theorem 1.1] for each $n\in N$. When $h$ is replaced by $h_{n}$, we also get the weak uniqueness of solutions by applying the results from Section 8. Furthermore, we can apply the duality relation from Corollary 6.15 for $X^{n}$, which is conditional on its existence.
@@ 7.1 Tightness
The proof follows from standard estimates as in the proof of [35, Theorem 2.6]. ∎
Lemma 7.1 demonstrates by [35, Lemma 6.3] that the sequence $(X^{n})_{n\geq 1}$ of $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}(\mathbb{R}))$ valued random variables is tight. Thus, by Prohorov’s theorem and Skorohod’s representation theorem, there exists a sub-sequence $(X^{n_{k}})_{k\geq 1}$ and another sequence $(\tilde{X}^{n_{k}})_{k\geq 1}$ of $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}(\mathbb{R}))$ valued random variables such that the following holds.
@@ 7.2 Convergence of $h_{n}(X^{n})$
The proof is similar to the proof of [4, Lemma 5.3]. First, we prove the $L^{1}$ convergence. Let $T_{0}>0$ be such that Proposition 4.3 holds for $X^{n}$, for all $n\geq 0$. Note that it is possible to choose $T_{0}>0$ uniformly in $n$, since it only depends on $\sup_{n}\nu_{n}^{1}(0,\infty)+\nu_{n}^{2}(0,\infty)$, where $\nu_{n}^{1},\nu_{n}^{2}$ are given by
$\nu_{n}^{i}=\nu^{i}_{|[0,n]}+d_{i}\delta_{n},\quad i\in\{1,2\},n\geq 1.$
For simplicity denote $X_{t}(x)$ by $Z$ and $X^{n}_{t}(x)$ by $Z_{n}$. Below we use the fact that almost surely $Z_{n},Z\geq 0$. Thus we get
@@ 8 Proof of Weak Uniqueness
Since $f$ was arbitrary it follows by [19, Corollary 2.3] that $X^{1}_{t}\overset{d}{=}X_{t}^{2}$. ∎
@@
It is enough to prove weak uniqueness on the interval $0,T_{0}]$ and that $\mathbb{P}\circ(X_{T_{0}})^{-1}\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ and that $X_{T_{0}}$ satisfies (1.5), since $T_{0}$ does not depend on the initial value $X_{0}$. This follows along similar lines as Lemma 7.1 and by Lemma 6.11. Thus, it remains to prove uniqueness on $0,T_{0}]$. However, by using Lemma 8.1, this follows along the same lines as the proof of [15, Theorem 4.2]. It is only necessary to show that t $\mathbb{P}\circ(X_{t})^{-1}\in\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ and that $X_{t}$ satisfies (1.5) for all $t\in[0,T_{0}]$. However, this once again follows along similar lines as Lemma 7.1 and from Lemma 6.11. ∎
@@
We only prove the statement for the case $b_{1}\geq 0$. When $b_{1}<0$ the statement follows by a comparison with the solution when $b_{1}=0$, by using weak uniqueness, [35, Corollary 2.4] and standard approximation and tightness arguments. Thus, we assume without loss of generality that $b_{1}\geq 0$ from now on.
The proof relies on the duality relation. To this end, let $Y$ be the process constructed in the beginning of Section 6 with $Y_{0}=\infty\delta_{0}$ and let $I_{t}(Y_{0})$ for $t\geq 0$ be defined as in (6.8). By Lemma 6.2, there exists $T_{0}>0$ such that
$\mathbb{E}[\operatorname{e}^{|I_{T_{0}}(Y_{0})||}]<\infty.$ (9.1)
For $t=0$ the assertion (1.7) follows by the assumptions of the theorem. Thus, let $s\in(0,T_{0}]$. Using Proposition 4.3 and Remark 4.4 it follows that
@@
where $\bar{T}_{n,k}^{x}$ is time of the $k$-th jump of the process $Y(n\delta_{x})$. That is, $K^{n}_{t}$ is the number of jumps of $Y(n\delta_{x})$ on the time interval $(0,t]$. We furthermore denote $T_{n,k}^{x}=\bar{T}_{n,k}^{x}-\bar{T}_{n,k-1}^{x}$, for $k\geq 1$. We obtain the following decomposition
@@
Finally, we estimate $I_{2}^{n}(s,x)$. To this end, we break up the event into several parts that we can handle more easily.
@@
$\begin{split}&\mathbb{P}(U_{n,j}^{x}>R_{x}+R,j\in\{1,\ldots k-1\},U_{n,k}^{x}\leq R_{x}+R,K_{s}^{n}=m)\\ &\leq\mathbb{P}(U_{n,k}^{x}\leq R_{x}+R,U_{n,j}^{x}>x-j\eta_{x}\text{ for all }j\in\{1,\ldots k-1\},K_{s}^{n}=m)\\ &\qquad+\mathbb{P}(U_{n,j}^{x}\leq x-j\eta_{x}\text{ for some }j\in\{1,\ldots k-1\},K_{s}^{n}=m).\end{split}$ (9.9)
@@
where $\mathcal{G}_{n,x,k}$ is the $\sigma$-algebra generated by $(U_{n,1}^{x},\ldots U_{n,k-1}^{x},T_{n,1}^{x},\ldots T_{n,k}^{x})$ . For $y\in\mathbb{R}$, using Lemma 3.3, Proposition 3.6 and the construction of the process $Y(n\delta_{x})$ we get that
@@
where we used that $\eta_{x}\leq R_{x}$ for $x\geq R^{\ast}$. This proves (1.7). ∎
10 Proof of Theorem 1.4
The proof of Theorem 1.4 relies on a rescaling of (1.1) as well as a comparison with solutions to
$\begin{split}d_{t}\bar{X}^{\theta}_{t}(x)&=\frac{1}{6}\Delta\bar{X}^{\theta}_{t}(x)+\theta\bar{X}^{\theta}_{t}(x)-\big(\bar{X}^{\theta}_{t}(x)\big)^{2}+\sqrt{\bar{X}^{\theta}_{t}(x)}\dot{W}(t,x),\quad t>0,\quad x\in\mathbb{R},\\ \bar{X}^{\theta}_{0}(x)&=f(x),\quad x\in\mathbb{R},\end{split}$ (10.1)
where $\theta>0$ and $f\in\mathcal{C}_{c}^{+}$.
Lemma 10.1.
Let $a,b,c>0$ and let $X$ be a solution to (1.1). Then
$\tilde{X}_{t}(x)=cX_{at}(bx),\quad t\geq 0,\quad x\in\mathbb{R},$
is a solution to
$\begin{split}d_{t}\tilde{X}_{t}(x)&=\frac{a}{2b^{2}}\Delta\tilde{X}_{t}(x)+ach_{c}(\tilde{X}_{t}(x))+\frac{c^{1/2}a^{1/2}}{b^{1/2}}\sqrt{\tilde{X}_{t}(x)}\dot{W}(t,x),\quad x\in\mathbb{R},t\geq 0,\\ \tilde{X}_{0}(x)&=cX_{0}(bx),\quad x\in\mathbb{R},\end{split}$
where $h_{c}(x)=h(c^{-1}x)$ for $x\in\mathbb{R}$.
Proof.
This follows along the same lines as the proof of [28, Lemma 2.1.2]. ∎
Lemma 10.2.
Let
$h(x)=\theta(1-e^{-\lambda x})$
for some $\lambda,\theta>0$ and let $X$ be a solution to (1.1) with $X_{0}=f\in\mathcal{C}_{c}^{+}$ such that $f\neq 0$. Then it follows that
$\mathbb{P}(\langle X_{t},1\rangle>0,\forall t>0)>0.$
Proof.
By [28, Theorem 1] there exists $\theta^{\ast}>0$ such that
$\mathbb{P}(\langle\bar{X}^{\theta^{\ast}}_{t},1\rangle>0,\forall t>0)>0.$
Now, we set $a=1/3b^{2}$, and $c=3b^{-1}$ for $b>0$. Then, by Lemma 10.1, $\tilde{X}_{t}(x)=cX_{at}(bx)$ solves
$d_{t}\tilde{X}_{t}=\frac{1}{6}\Delta\tilde{X}_{t}(x)+b\theta(1-\operatorname{e}^{-\frac{b}{3}\lambda\tilde{X}_{t}(x)})+\sqrt{\tilde{X}_{t}(x)}\dot{W}(t,x),\quad x\in\mathbb{R},t\geq 0.$ (10.2)
Suppose, that there exists $b>0$ such that
$b\theta(1-\operatorname{e}^{-\frac{b}{3}\lambda x})\geq\theta^{\ast}x-x^{2},\quad\forall x\geq 0.$ (10.3)
By [22, Theorem 3.3] and standard approximation and tightness arguments it follows by weak uniqueness that $\tilde{X}$ and $\bar{X}$ can be defined on a common probability space such that for all $t\geq 0$ we have
$\tilde{X}_{t}(x)\geq\bar{X}^{\theta^{\ast}}_{t}(x),\quad x\in\mathbb{R},$
almost surely. Again by weak uniqueness and by continuity of $\tilde{X}$ and $\bar{X}$ we get
$\mathbb{P}(\tilde{X}_{t}(x)\geq\bar{X}^{\theta^{\ast}}_{t}(x),\forall t\geq 0,x\in\mathbb{R})=1.$
This in particular implies that $\mathbb{P}\langle(\tilde{X}_{t},1\rangle>0,\forall t>0)>0$ and thus $\mathbb{P}(\langle X_{t},1\rangle>0,\forall t>0)>0$. Thus, it only remains to establish (10.3), which is equivalent to proving
$\frac{b^{3}\lambda^{2}}{9}\theta(1-\operatorname{e}^{-x})\geq\theta^{\ast}\frac{b\lambda}{3}x-x^{2},\quad\forall x\geq 0,$
for some $b>0$. For $x\in[0,1]$ we may use that $(1-\operatorname{e}^{-x})\geq\frac{1}{2}x$ and thus in this regime it is sufficient to choose $b>0$ sufficiently large so that
$\frac{b^{3}\lambda^{2}}{9}\theta\frac{1}{2}\geq\theta^{\ast}\frac{b\lambda}{3}.$
Furthermore, for note that
$\Big(\theta^{\ast}\frac{b\lambda}{3}\Big)^{2}\frac{1}{4}\geq\theta^{\ast}\frac{b\lambda}{3}x-x^{2},$
for $x\in\mathbb{R}$. Thus, for $x>1$, we need to choose $b>0$ sufficiently large such that also
$\frac{b^{3}\lambda^{2}}{9}\theta(1-\operatorname{e}^{-1})\geq\Big(\theta^{\ast}\frac{b\lambda}{3}\Big)^{2}\frac{1}{4}.$
That means, (10.3) is satisfied for
$b>\max\Big(\sqrt{\frac{6\theta^{\ast}}{\lambda\theta}},\frac{(\theta^{\ast})^{2}}{4\theta(1-e^{-1})}\Big),$
which finishes the proof. ∎
Proof of Theorem 1.4.
Suppose that there exists $\lambda,\theta>0$ such that
$h(x)\geq\theta(1-\operatorname{e}^{-\lambda x}),\quad\forall x\geq 0.$ (10.4)
Then, by weak uniqueness, [35, Corollary 2.4] and standard approximation and tightness arguments we have that
$\mathbb{P}(\langle X_{t},1\rangle>0,\forall t>0)>\mathbb{P}(\langle\hat{X}_{t},1\rangle>0,\forall t>0)$
where $\hat{X}$ is the solution to (1.1) with $h(x)=\theta(1-\operatorname{e}^{-\lambda x})$ for all $x\in\mathbb{R}$. Thus, by Lemma 10.2 it is sufficient to verify (10.4). If $b_{1}>0$, then (10.4) is clearly satisfied for all $\lambda>0$ and $\theta\leq b_{1}$. When $b_{1}=0$ we have the estimate
$\int_{0}^{\infty}(1-\operatorname{e}^{-\tilde{\lambda}x})\nu^{1}(d\tilde{\lambda})\geq(1-\operatorname{e}^{-\lambda x})\nu^{1}([\lambda,\infty)),$
where $\lambda>0$ is chosen in such a way that $\nu^{1}([\lambda,\infty))>0$. Then, (10.4) is clearly satisfied for $\theta\leq\nu^{1}([\lambda,\infty))$. ∎